# Cycle Matroids of Graphings: From Convergence to Duality

### v2

@@ 5.2 A matroid attaining the cost
As we are in the non-hyperfinite case, the dual $\mathcal{G}^{*}$ has no $2$-ended components $\mu^{*}$-almost surely. By [7], one can pick a 1-ended hyperfinite spanning forest $F^{*}\subseteq\mathcal{G}^{*}$. The corresponding measure $\alpha_{F^{*}}$ is in $\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$, implying $\tilde{\mu}-\alpha_{F^{*}}\in\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$. This measure, considered on $E(\mathcal{G})$, corresponds to $E\setminus F$. The key idea, also exploited in [7], is that $F^{*}$ being $1$-ended almost everywhere means that $E\setminus F$ is connected on almost every component of $\mathcal{G}$. That is, $E\setminus F$ is a treeing of $\mathcal{G}$. Therefore, $\rho^{D}_{\mathcal{G}^{*}}(E)=\alpha_{E\setminus F}(E)=\tilde{\mu}(E\setminus F)=\textrm{cost}(\mathcal{G})$, where the last inequality holds by Theorem 2.1. ∎
@@ 5.3 Free versus wired spanning forests
Let $\mathcal{G}$ be a graphing with a Borel order $<$ on the edges. We have already seen in Remark 4.10 that the algorithm for finding bases with respect to $\rho_{\mathcal{G}}$ produces the wired maximal spanning forest, denoted by $\mathcal{T}_{w}$. There is a natural counterpart, the free maximal spanning forest. We again refer to [20, Chapter 11] for a thorough overview, with the caveat that here we treat arbitrary Borel orders on the edge set.
We define the free maximal spanning forest $\mathcal{T}_{f}$ by simultaneously deleting all edges $e$ of $\mathcal{G}$ for which there exists a cycle $C$ containing $e$ in which $e$ is $<$-minimal. It is straightforward to check that $\mathcal{T}_{f}$ is acyclic, and all of its components are infinite. Moreover, $\mathcal{T}_{w}\subseteq\mathcal{T}_{f}$.
@@
When $\mathcal{G}$ and $\mathcal{G}^{*}$ are dual planar graphings and the Borel total ordering is the same on $E$ and $E^{*}$, the free maximal spanning forest of $\mathcal{G}$ is exactly the complement of the wired minimal spanning forest of $\mathcal{G}^{*}$. ∎
@@
Based on Theorem 1.10, one could hope to use the free maximal spanning forest for arbitrary, not necessarily planar graphings to define an interesting matroid that goes beyond hyperfinite spanning forests. The problem is that the free maximal spanning forest built on two different Borel orders might have different edge measure, and so they cannot form bases of the same matroid. For example, let $\mathcal{G}$ be the Schreier graphing of a free action of $F_{2}\times\mathbb{Z}$. Let $<_{1}$ be such that all $F_{2}$-edges are larger than all $\mathbb{Z}$-edges, and the ordering is reversed for $<_{2}$. Then, $\mathcal{T}^{<_{1}}_{f}$ contains exactly the $F_{2}$-edges while $\mathcal{T}^{<_{2}}_{f}$ contains exactly the $\mathbb{Z}$-edges, so their edge measures are $2$ and $1$, respectively.

### v3

@@ 5.2 A matroid attaining the cost
As we are in the non-hyperfinite case, the dual $\mathbf{G}^{*}$ has no $2$-ended components $\mu^{*}$-almost surely. By [10], one can pick a 1-ended hyperfinite spanning forest $F^{*}\subseteq E^{*}$. The corresponding measure $\alpha_{F^{*}}$ is in $\text{\sf bmm}^{+}(\rho_{\mathbf{G}^{*}})$, implying $\eta_{\mathbf{G}^{*}}-\alpha_{F^{*}}\in\text{\sf bmm}^{+}(\rho^{*}_{\mathbf{G}^{*}})$. This measure, considered on $E(\mathbf{G})$, corresponds to $E\setminus F$. The key idea, also exploited in [10], is that $F^{*}$ being $1$-ended almost everywhere means that $E\setminus F$ is connected on almost every component of $\mathbf{G}$. That is, $E\setminus F$ is a treeing of $\mathbf{G}$. Therefore, $\rho^{*}_{\mathbf{G}^{*}}(E)=\alpha_{E\setminus F}(E)=\eta_{\mathbf{G}}(E\setminus F)=\textrm{\rm cost}(\mathbf{G})$, where the last inequality holds by Theorem 2.4. ∎
@@ 5.3 Free versus wired spanning forests
Let $\mathbf{G}$ be a graphing with a Borel order $<$ on the edges. We have already seen in Remark 4.9 that the algorithm for finding bases with respect to $\rho_{\mathbf{G}}$ produces the wired maximal spanning forest, denoted by $\mathbf{T}_{w}$. Recall that the free maximal spanning forest is a natural counterpart. We again refer to [24, Chapter 11] for a thorough overview, with the caveat that here we treat arbitrary Borel orders on the edge set.
@@
When $\mathbf{G}$ and $\mathbf{G}^{*}$ are dual planar graphings and the Borel total ordering is the same on $E$ and $E^{*}$, the free maximal spanning forest of $\mathbf{G}$ is exactly the complement of the wired minimal spanning forest of $\mathbf{G}^{*}$. ∎
@@
Based on Theorem 1.8, one could hope to use the free maximal spanning forest for arbitrary, not necessarily planar graphings to define an interesting matroid that goes beyond hyperfinite spanning forests. The problem is that the free maximal spanning forest built on two different Borel orders might have different edge measure, and so they cannot form bases of the same matroid. For example, let $\mathbf{G}$ be the Schreier graphing of a free action of $F_{2}\times\mathbb{Z}$. Let $<_{1}$ be such that all $F_{2}$-edges are larger than all $\mathbb{Z}$-edges, and the ordering is reversed for $<_{2}$. Then, $\mathbf{T}^{<_{1}}_{f}$ contains exactly the $F_{2}$-edges while $\mathbf{T}^{<_{2}}_{f}$ contains exactly the $\mathbb{Z}$-edges, so their edge measures are $2$ and $1$, respectively.